### v2

# Anisotropy of quadratic forms over global fields of characteristic $\neq$ 2 is Diophantine

Abstract.
We prove that the set of anisotropic quadratic forms over global fields of characteristic different from 2 is a diophantine set. Our proof builds upon and extends the method of Koenigsmann, using tools from class field theory, local-global principles, and recent advances on the diophantine definability of non-norm sets over global fields.
1. Introduction
Let $K$ be a global field. Roughly speaking, a diophantine set is a set that can be described by the solvability of a family of polynomial equations. Equivalently, we may give the following definition.
Definition 1.1.
A subset $S\subset K^{n}$ is called a diophantine set over $K$ if there exists a polynomial
$f(a_{1},\ldots,a_{n};x_{1},\ldots,x_{m})\in K[a_{1},\ldots,a_{n};x_{1},\ldots,x_{m}]$
such that for each $(a_{1},\ldots,a_{n})\in K^{n}$ we have
$(a_{1},\ldots,a_{n})\in S\iff\exists(x_{1},\ldots,x_{m})\in K^{m}\text{ such that }f(a_{1},\ldots,a_{n};x_{1},\ldots,x_{m})=0.$
The study of diophantine sets is closely connected with Hilbert’s Tenth Problem. Matiyasevich proved that the problem admits a negative answer over $\mathbb{Z}$ (see [Mat70]). Furthermore, if $\mathbb{Z}$ is diophantine over $\mathbb{Q}$, then Hilbert’s Tenth Problem over $\mathbb{Q}$ would also have a negative answer. However, this remains an open problem.
Besides their connection to foundational questions in logic and number theory, the study of diophantine sets is of intrinsic interest as an inverse to the problem of solving polynomial equations. While the classical problem asks for the solutions of given polynomial equations, the study of diophantine sets addresses the reverse question: which subsets of $K^{n}$ can be realized as the set of solutions to some family of polynomial equations.
A significant advance in this direction was achieved by Poonen [Poo09], who proved, using the Brauer–Manin obstruction, that the set of non-squares in number fields is diophantine. This provided one of the first nontrivial examples of such sets.
This result has been generalized to various situations in recent years. Colliot-Thélène and Van Geel [CTG15] extended Poonen’s result to the complements of $n$-th powers over global fields. Koenigsmann [Koe16] introduced powerful new methods for constructing diophantine sets over $\mathbb{Q}$, notably the set of non-norms for quadratic extensions.
More generally, for a quadratic extension $L/K$ of a field $K$, the set of non-norms
$\{x\in K^{\times}\mid x\notin N_{L/K}(L^{\times})\}$
is diophantine. In fact, a stronger statement holds:
$\{(x,y)\in(K^{\times})^{2}\mid x\notin N_{K(\sqrt{y})/K}(K(\sqrt{y})^{\times})\}$
is diophantine. Park [Par13] extended Koenigsmann’s method to number fields, and Eisenträger and Morrison [EM18] later applied it to global function fields of odd characteristic. Morrison [Mor19] further generalized the result to cyclic extensions of degree prime to the characteristic.
Another result was obtained by Philip Dittmann [Dit18]. He generalized the method uniformly to apply to any global field and proved that the set of coefficients of univariate polynomials that have no solutions in a global field is diophantine. As a corollary, he showed that the set of coefficients of irreducible multivariate polynomials is diophantine.
In this paper, we turn to diophantine sets defined by quadratic forms. Our goal is to show that the set of coefficients of anisotropic quadratic forms over a global field of characteristic not equal to $2$ is a diophantine set. In any case, the set of coefficients of isotropic quadratic forms is obviously diophantine. So, the set we deal with here is the complement of this obvious diophantine set, just like the works we mentioned above. Our main theorem is the following.
Theorem 1.2.
Let $K$ be a global field with $\textup{char}(K)\neq 2$ and let $m\in\mathbb{N}$. Then
$\bigl\{(a_{i})\in(K^{\times})^{m}:\sum_{i=1}^{m}a_{i}x_{i}^{2}\text{ is anisotropic over }K\bigr\}$
forms a diophantine set over $K$.
Remark 1.3.
In the notation $(K^{\times})^{m}$ we mean the set of $m$-tuples of elements of $K^{\times}$, not the set of $m$th powers in $K^{\times}$, which we denote by ${K^{\times}}^{m}$.
We say that a quadratic form $f$ represents 0 over $K$ if there exists a non-trivial tuple $(x_{1},\dots,x_{m})\in K^{m}\setminus\{0\}$ such that $f(x_{1},\dots,x_{m})=0$.
As a corollary, we obtain the following:
Corollary 1.4.
Let $K$ be a global field with $\textup{char}(K)\neq 2$ and let $m\in\mathbb{N}$. Then
$\bigl\{(a_{ij})\in(K^{\times})^{m^{2}}:\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}\text{ is anisotropic over }K\bigr\}$
forms a diophantine set over $K$.
To prove Theorem 1.2, we consider different values of $m$. For $m\geq 5$, the result follows from local-global principles. For $m=1,2,$ or $3$, the result is already known in the context of “non-squares” and “non-norms”. The only genuinely new case is $m=4$, for which we employ generalizations of Koenigsmann’s method as developed in [Dit18, Par13, EM18].
2. Preliminaries
From now on, we use $S_{K}$ to denote the set of primes of $K$, including both finite and infinite primes. By a prime of $K$ we mean an equivalence class of nontrivial absolute values of $K$.
For a number field $K$, we write $S_{K}^{\infty}$ for the set of infinite primes. For $\mathfrak{p}\in S_{K}^{\infty}$, if the corresponding completion $K_{\mathfrak{p}}$ is isomorphic to the real field $\mathbb{R}$, we call $\mathfrak{p}$ a real infinite prime; if $K_{\mathfrak{p}}\simeq\mathbb{C}$, then $\mathfrak{p}$ is called a complex infinite prime.
[…]
For $a,b\in K_{\mathfrak{p}}^{\times}$, the notation $(a,b)_{\mathfrak{p}}$ denotes the Hilbert symbol at $\mathfrak{p}$. Note that the Hilbert symbol $(a,b)_{\mathfrak{p}}$ is a nondegenerate bilinear pairing on $K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$. For $a\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$, we define the homomorphism of groups
$T_{a}:K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}\rightarrow\{\pm 1\}$
by
$x\mapsto(x,a)_{\mathfrak{p}},$
where we regard $\{\pm 1\}$ as a group by multiplication.
Lemma 2.1.
With the notations above, for $\epsilon,\epsilon^{\prime}\in\{\pm 1\}$ and nontrivial $a,a^{\prime}\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$, we have $T_{a}^{-1}(\epsilon)\cap T^{-1}_{a^{\prime}}(\epsilon^{\prime})=\emptyset\Leftrightarrow a=a^{\prime}\wedge\epsilon=-\epsilon^{\prime}$.
Proof.
$\Leftarrow:$ This is obvious.
$\Rightarrow:$ By 63:9 of [O’M00] we know that the index $K_{\mathfrak{p}}^{\times}:{K_{\mathfrak{p}}^{\times}}^{2}]$ is finite. $T_{a}$ is nontrivial since $a\neq 1$ and the Hilbert symbol is nondegenerate, hence the cardinality of $T_{a}^{-1}(1)$ and $T_{a}^{-1}(-1)$ are both half of $K_{\mathfrak{p}}^{\times}:{K_{\mathfrak{p}}^{\times}}^{2}]$.
For $\epsilon,\epsilon^{\prime}\in\{\pm 1\}$ and nontrivial $a,a^{\prime}\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$, we have
$T_{a}^{-1}(\epsilon)\cap T^{-1}_{a^{\prime}}(\epsilon^{\prime})=\emptyset\Leftrightarrow T_{a}^{-1}(\epsilon)\sqcup T^{-1}_{a^{\prime}}(\epsilon^{\prime})=K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$
by the reason of cardinality.
The complement of $T_{a}^{-1}(1)$ in $K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$ is $T_{a}^{-1}(-1)$, so
$T_{a}^{-1}(\epsilon)\sqcup T^{-1}_{a^{\prime}}(\epsilon^{\prime})=K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$
is equivalent to
[…]
However, $T_{a}^{-1}(1)=T_{a^{\prime}}^{-1}(-1)$ is impossible, since this will imply
$(x,aa^{\prime})_{\mathfrak{p}}=(x,a)_{\mathfrak{p}}(x,a^{\prime})_{\mathfrak{p}}=-1,\forall x\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2},$
which cannot happen.
Also, it is easy to verify that
$T_{a}^{-1}(1)=T_{a^{\prime}}^{-1}(1)$ $\Leftrightarrow(x,a)_{\mathfrak{p}}=(x,a^{\prime})_{\mathfrak{p}},\forall x\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$
$\Leftrightarrow(x,aa^{\prime})_{\mathfrak{p}}=1,\forall x\in K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$
$\Leftrightarrow T_{aa^{\prime}}\textup{ is trivial}$
$\Leftrightarrow aa^{\prime}=1$
$\Leftrightarrow a=a^{\prime}.$
In conclusion, we have $T_{a}^{-1}(\epsilon)\cap T^{-1}_{a^{\prime}}(\epsilon^{\prime})=\emptyset\Leftrightarrow a=a^{\prime}\wedge\epsilon=-\epsilon^{\prime}$. ∎
2.1. Quadratic forms
The arithmetic of quadratic forms over a global field is well understood. In the case of fields with characteristic not equal to $2$, the situation is even simpler, as the local-global principle applies.
For proving the Theorem 2.3 below, we need the following lemma:
Lemma 2.2.
For $(a_{1},a_{2},a_{3},a_{4})\in(K_{\mathfrak{p}}^{\times})^{4}$, The following two conditions are equivalent:
  • (1)
[…]
  • (2)
[…]
    $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}=-a_{3}x_{3}^{2}-a_{4}x_{4}^{2}=x^{\prime}$
Proof.
[…]
For (1) $\Rightarrow$ (2), suppose that there exists
[…]
such that
$a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2}=0.$
If $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}\neq 0$, take $x^{\prime}=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ then we are done. If $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}=0$, then the quadratic form $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ is isotropic, hence universal by 42:10 of [O’M00]. Then we can take $x^{\prime}=-a_{3}x_{3}^{2}-a_{4}x_{4}^{2}$ for arbitrary $(x_{3},x_{4})\in(K_{\mathfrak{p}})^{2}$ such that $x^{\prime}\neq 0$. ∎
Theorem 2.3.
Let $K$ be a global field with $\textup{char}(K)\neq 2$, and let $f$ be a non-degenerate quadratic form in $m$ variables over $K$. Then the following statements hold:
  • $f$ is isotropic over $K$ if and only if $f$ is isotropic over $K_{\mathfrak{p}}$ for every prime $\mathfrak{p}\in S_{K}$.
  • If $m\geq 5$ and $\mathfrak{p}\in S_{K}$ is a non-archimedean place, then $f$ is isotropic over $K_{\mathfrak{p}}$.
  • If $m=4$, $\mathfrak{p}\in S_{K}$ is non-archimedean, and
    $f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2},$
    then $f$ is anisotropic over $K_{\mathfrak{p}}$ if and only if $a_{1}a_{2}a_{3}a_{4}\in{K_{\mathfrak{p}}}^{\times 2}$ and $(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}$.
Proof.
The first item follows from Theorem 66:1 in [O’M00], and the second item from Theorem 63:19 in [O’M00]. The third item is stated in Theorem 6 of Chapter IV in [Ser73] for $K=\mathbb{Q}$. The same proof can be naturally generalized to global fields. We provide a detailed proof of this generalization below, following the method in [Ser73].
By Lemma 2.2, $f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2}$ is isotropic over $K_{\mathfrak{p}}$ if and only if there exists a nonerzo $x^{\prime}\in K_{\mathfrak{p}}$ such that the two forms $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ and $a_{3}x_{3}^{2}-a_{4}x_{4}^{2}$ represents $x^{\prime}$. Let $x$ denote the coset of $x^{\prime}$ in $K_{\mathfrak{p}}^{\times}/{K_{\mathfrak{p}}^{\times}}^{2}$. The form $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ represents $x^{\prime}$ is equivalent to $(xa_{1},xa_{2})_{\mathfrak{p}}=1$. After a calculation of Hilbert symbol, we get:
$(xa_{1},xa_{2})_{\mathfrak{p}}$ $=(xa_{1},x)_{\mathfrak{p}}(xa_{1},a_{2})_{\mathfrak{p}}$
$=(x,x)_{\mathfrak{p}}(a_{1},x)_{\mathfrak{p}}(x,a_{2})_{\mathfrak{p}}(a_{1},a_{2})_{\mathfrak{p}}$
$=(x,-1)_{\mathfrak{p}}(x,a_{1})_{\mathfrak{p}}(x,a_{2})_{\mathfrak{p}}(a_{1},a_{2})_{\mathfrak{p}}$
$=(x,-a_{1}a_{2})_{\mathfrak{p}}(a_{1},a_{2})_{\mathfrak{p}}.$
So $a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ represents $x^{\prime}$ is equivalent to
$(x,-a_{1}a_{2})_{\mathfrak{p}}=(a_{1},a_{2})_{\mathfrak{p}},$
which means exactly $x\in T_{-a_{1}a_{2}}^{-1}((a_{1},a_{2})_{\mathfrak{p}})$.
For the same reason, $a_{3}x_{1}^{2}-a_{4}x_{2}^{2}$ represents $x^{\prime}$ is equivalent to
$(x,-a_{3}a_{4})_{\mathfrak{p}}=(-a_{3},-a_{4})_{\mathfrak{p}},$
which means exactly $x\in T_{-a_{3}a_{4}}^{-1}((-a_{3},-a_{4})_{\mathfrak{p}})$.
Hence the absence of such an $x$ is equivalent to $T_{-a_{1}a_{2}}^{-1}((a_{1},a_{2})_{\mathfrak{p}})\cap T_{-a_{3}a_{4}}^{-1}((-a_{3},-a_{4})_{\mathfrak{p}})=\emptyset$. By Lemma 2.1, this is equivalent to
[…]
2.2. Global class field theory
In this section, we recall some standard results from global class field theory. We begin by introducing the notion of a modulus, which generalizes the integer $m$ in classical modular arithmetic.
Definition 2.4.
Let $K$ be a global field and $\mathcal{O}_{K}$ its ring of integers. A modulus of $K$ is a formal product of primes of $K$
$\mathfrak{m}=\prod_{\mathfrak{p}}\mathfrak{p}^{\mathfrak{m}(\mathfrak{p})},$
where each exponent $\mathfrak{m}(\mathfrak{p})\geq 0$, and the following conditions are satisfied:
  • $\mathfrak{m}(\mathfrak{p})=0$ for all but finitely many primes $\mathfrak{p}$;
  • $\mathfrak{m}(\mathfrak{p})=0$ for every infinite complex prime $\mathfrak{p}$;
  • $\mathfrak{m}(\mathfrak{p})\in\{0,1\}$ for every infinite real prime $\mathfrak{p}$.
In the number field case, a modulus can be written as
$\mathfrak{m}=\mathfrak{m}_{0}\cdot\mathfrak{m}_{\infty},$
where $\mathfrak{m}_{0}$ denotes the finite part and $\mathfrak{m}_{\infty}$ the infinite part. The notion of congruence extends naturally to this setting.
Definition 2.5.
Let $a,b\in K^{\times}$, and let $\mathfrak{m}$ be a modulus of $K$. For a finite or real infinite prime $\mathfrak{p}$, we define
$a\equiv^{*}b\pmod{\mathfrak{p}^{\mathfrak{m}(\mathfrak{p})}}$
as follows:
  • If $\mathfrak{p}$ is finite, then
    $a\equiv^{*}b\pmod{\mathfrak{p}^{\mathfrak{m}(\mathfrak{p})}}\quad\Longleftrightarrow\quad v_{\mathfrak{p}}\!\left(\tfrac{a}{b}-1\right)\geq\mathfrak{m}(\mathfrak{p}).$
  • If $\mathfrak{p}$ is real, then
    $a\equiv^{*}b\pmod{\mathfrak{p}}\quad\Longleftrightarrow\quad\frac{a}{b}>0$
    under the real embedding associated with $\mathfrak{p}$.
Moreover, we write
$a\equiv^{*}b\pmod{\mathfrak{m}}$
if the above holds for every $\mathfrak{p}$ with $\mathfrak{m}(\mathfrak{p})>0$.
Let $I_{K}$ denote the group of fractional ideals of $K$, and let
$I_{\mathfrak{m}}:=\langle\mathfrak{p}\subseteq\mathcal{O}_{K}:\mathfrak{m}(\mathfrak{p})=0\rangle$
be the subgroup of $I_{K}$ generated by all prime ideals $\mathfrak{p}$ of $K$ that do not divide $\mathfrak{m}$.
Definition 2.6.
The ray modulo $\mathfrak{m}$ is
$K_{\mathfrak{m},1}=\{a\in K^{\times}:a\equiv^{*}1\pmod{\mathfrak{m}}\}$.
Every $\alpha\in K^{\times}$ defines a principal fractional ideal
$(\alpha):=\alpha\mathcal{O}_{K}\;\in\;I_{K}.$
If $\alpha\in K_{\mathfrak{m},1}$, then $(\alpha)\in I_{\mathfrak{m}}$, so we obtain a well-defined homomorphism
$i:K_{\mathfrak{m},1}\;\longrightarrow\;I_{\mathfrak{m}},\qquad\alpha\longmapsto(\alpha).$
Definition 2.7.
Let
$P_{\mathfrak{m}}:=i(K_{\mathfrak{m},1})\;\leq\;I_{\mathfrak{m}}.$
The ray class group modulo $\mathfrak{m}$ is defined as
$C_{\mathfrak{m}}:=I_{\mathfrak{m}}/P_{\mathfrak{m}}.$
Let $L/K$ be a finite abelian extension. For a prime ideal $\mathfrak{p}$ of $K$ unramified in $L$, denote by
$(\mathfrak{p},L/K)\;\in\;\mathrm{Gal}(L/K)$
the associated Frobenius automorphism.
If $\mathfrak{m}$ is a modulus of $K$ divisible by all primes that ramify in $L$, the global Artin map is the homomorphism
$\psi_{L/K}:I_{\mathfrak{m}}\;\longrightarrow\;\mathrm{Gal}(L/K),$
defined on a fractional ideal
$\mathfrak{a}\;=\;\prod\mathfrak{p}_{i}^{e_{i}}\;\in\;I_{\mathfrak{m}}$
by
$\psi_{L/K}(\mathfrak{a})\;:=\;\prod(\mathfrak{p}_{i},L/K)^{e_{i}}.$
Theorem 2.8 (Artin Reciprocity).
Let $L/K$ be a finite abelian extension. There exists a modulus $\mathfrak{m}$ of $K$, containing all primes of $K$ that ramify in $L$, such that the global Artin map
$\psi_{L/K}:I_{\mathfrak{m}}\longrightarrow\mathrm{Gal}(L/K)$
is surjective, with kernel
$\ker(\psi_{L/K})\;=\;P_{\mathfrak{m}}\cdot N_{L/K}\big(I_{L}(\mathfrak{m}^{\prime})\big),$
where $\mathfrak{m}^{\prime}$ is the modulus of $L$ associated to $\mathfrak{m}$, defined as follows.
If
$\mathfrak{m}\;=\;\prod_{\mathfrak{p}}\mathfrak{p}^{n_{\mathfrak{p}}}\cdot\prod_{\begin{subarray}{c}\omega\text{ real}\end{subarray}}\omega^{n_{\omega}},\qquad n_{\omega}\in\{0,1\},$
then
$\mathfrak{m}^{\prime}\;=\;\prod_{\mathfrak{P}\mid\mathfrak{p}}\mathfrak{P}^{n_{\mathfrak{p}}}\cdot\prod_{\begin{subarray}{c}\Omega\mid\omega\text{ real}\end{subarray}}\Omega^{n_{\omega}},$
where the first product runs over all finite primes $\mathfrak{P}$ of $L$ above $\mathfrak{p}$, and the second runs over all real embeddings $\Omega$ of $L$ extending the real embedding $\omega$ of $K$.
We call a modulus $\mathfrak{m}$ as in Theorem 2.8 an admissible modulus for the extension $L/K$.
Theorem 2.8 is stated for number fields and function fields in various standard references. For instance, it appears as Theorem 2.1 in [EM18] in the case of global function fields, and as Theorem 3.3 in [Par13] in the case of number fields. For a standard exposition of the Artin symbol, we refer to [Lan86], where a proof of Artin reciprocity is given in Theorems 1–3.
3. Main Method
In this section, we recall the methods used in [Koe16, Par13, EM18, Dit18] to construct diophantine sets. These methods will serve as the main ingredients in the proof of our main result.
3.1. Notations and Some Examples
Let $K$ be a field, and let $A$ be a finite-dimensional central simple algebra over $K$. We define
$S(A/K):=\{\textup{Trd}(x):x\in A,\ \textup{Nrd}(x)=1\},\qquad T(A/K):=S(A/K)-S(A/K),$
where Trd and Nrd denote the reduced trace and norm, respectively.
From now on, let $K$ be a global field of $\textup{char}(K)\neq 2$, and let $A$ be the quaternion algebra $H_{a,b}$ for $a,b\in K^{\times}$ defined by
$H_{a,b}:=K\oplus\alpha K\oplus\beta K\oplus\alpha\beta K,$
where $\alpha^{2}=a$, $\beta^{2}=b$, $\alpha\beta=-\beta\alpha$.
Note that for the quaternion algebra $H_{a,b}$ over a global field $K$, we have
T(A/K):=S(A/K)+S(A/K)=$S(A/K)$,T(A/K):=S(A/K)+S(A/K)=S(A/K)-S(A/K),
since for any $x\in A$ with $\mathrm{Nrd}(x)=1$, we have $x\in A$ and $\mathrm{Nrd}(-x)=\mathrm{Nrd}(x)=1$, and moreover $\mathrm{Trd}(-x)=-\mathrm{Trd}(x)$. Hence $x\in S(A/K)$ implies $x\in S(A/K)$, which ensures that S(A/K)+S(A/K)=S(A/K)S(A/K)S(A/K)+S(A/K)=S(A/K)-S(A/K).
For $\mathfrak{p}\in S_{K}^{f}$, let $R_{\mathfrak{p}}$ be its valuation ring in $K_{\mathfrak{p}}$, and define
$\mathcal{O}_{\mathfrak{p}}:=(\mathcal{O}_{K})_{\mathfrak{p}}=R_{\mathfrak{p}}\cap K$
as the local ring of $\mathfrak{p}$ in $K$. For simplicity, we write $S(A/K)$ as $S_{a,b}$ and $T(A/K)$ as $T_{a,b}$, since $A$ is the quaternion algebra as stated.
Lemma 3.1.
If $A/K$ splits at all real places of $K$ (which is always the case if $K$ is a global function field or $a,b\in K^{\times}$ are totally positive if $K$ is a number field), then
$T_{a,b}=\bigcap_{\mathfrak{q}\in\Delta_{a,b}}\mathcal{O}_{\mathfrak{q}},$
where $\Delta_{a,b}$ is the finite set of places of $K$ at which $A$ does not split.
Proof.
This is a special case of Proposition 2.9 in [Dit18]. ∎
Remark 3.2.
$T_{a,b}$ is a diophantine set over $K$ by definition, and this is the basis to prove that many other sets are diophantine in [EM18],[Dit18] and [Par13]. Actually, $T_{a,b}$ is uniformly existential defined in $a$ and $b$ since we have the following formula for $S_{a,b}$
$S_{a,b}=\{2x_{1}\in K:\exists x_{2},x_{3},x_{4}\textup{ with }x_{1}^{2}-ax_{1}^{2}-bx_{3}^{2}+abx_{4}^{2}=1\}$
as well as the definition $T_{a,b}=S_{a,b}+S_{a,b}$.
The following definition is given in Definition 9 of [Koe16] for $K=\mathbb{Q}$, and also given after the proof of Lemma 3.8 in [Par13] as well as after the proof of Lemma 3.3 in [EM18].
Definition 3.3.
Let $L:=K(\sqrt{a},\sqrt{b})/K$ be an abelian extension with Galois group isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{2}$ and $\psi_{L/K}:I_{\mathfrak{m}}\rightarrow\textup{Gal}(L/K)$ be the global Artin map with the admissible modulus $\mathfrak{m}$. We define the following notations:
  • $\mathbb{P}(p):=\{\mathfrak{p}\in S_{K}^{f}:v_{\mathfrak{p}}(p)\text{ is odd}\}$.
  • For $(i,j)\in\textup{Gal}(L/K)$, $i,j\in\{\pm 1\}$, we define $\mathbb{P}^{(i,j)}=\{\mathfrak{p}\in S_{K}^{f}:\mathfrak{p}\in I_{m}\text{ and }\psi_{L/K}(\mathfrak{p})=(i,j)\}$.
  • $\mathbb{P}^{(i,j)}(p):=\mathbb{P}(p)\cap\mathbb{P}^{(i,j)}$.
Here, $(1,1)$ is the identity in the Galois group. (1,1)(-1,1) is the element which sends $\sqrt{a}$ to $\sqrt{a}$ and fixes $\sqrt{b}$. (1,1)(1,-1) is the element which sends $\sqrt{b}$ to $\sqrt{b}$ and fixes $\sqrt{a}$. Finally, (1,1)(-1,-1) is the element sends $\sqrt{a}$ to $\sqrt{a}$ and $\sqrt{b}$ to $\sqrt{b}$.
The following lemma is a summary of some results in [Par13] and [EM18].
Lemma 3.4.
(1) If $K$ is a number field, then we can find $a,b\in K^{\times}$ with disjoint support and $\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)\cong(\mathbb{Z}/2\mathbb{Z})^{2}$ such that for any $p\in K^{\times}$ co-prime to the admissible modulus $\mathfrak{m}$, we have
  • $\mathbb{P}^{(-1,-1)}(p)=\Delta_{a,p}\cap\Delta_{b,p}$,
  • $\mathbb{P}^{(-1,1)}(p)=\Delta_{a,p}\cap\Delta_{ab,p}$,
  • $\mathbb{P}^{(1,-1)}(p)=\Delta_{b,p}\cap\Delta_{ab,p}$.
(2) If $K$ is a global function field with $\textup{char}(K)\neq 2$, then we can find $a,b,c,d\in K^{\times}$ with $\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)=(\mathbb{Z}/2\mathbb{Z})^{2}$ such that for any $p\in K^{\times}$ coprime to the the admissible modulus $\mathfrak{m}$, we have
  • $\mathbb{P}^{(-1,-1)}(p)=\Delta_{a,p}\cap\Delta_{b,p}$,
  • $\mathbb{P}^{(-1,1)}(p)=\Delta_{a,p}\cap\Delta_{ab,p}\cap\Delta_{a,cp}$,
  • $\mathbb{P}^{(1,-1)}(p)=\Delta_{b,p}\cap\Delta_{ab,p}\cap\Delta_{b,dp}$.
Proof.
First we explain (1), this is the case proved in [Par13]. From Lemma 3.9 of [Par13] we know that for any $p$ coprime to $\mathfrak{m}$, the sets at the left-hand and right-hand sides of each equation differ at most by the primes dividing the modulus. And then Lemma 3.19 and Lemma 3.20 in [Par13] give a carefully chosen $a$ and $b$ such that the two sets are equal.
The case (2) is Corollary 3.11 in [EM18].
The support of $a$ and $b$ are disjoint by Lemma 3.19 (4) of [Par13] and Lemma 3.8 (4) of [EM18]. ∎
From now on, unless otherwise stated, our parameters $a,b$ or $a,b,c,d$ are the same as in Lemma 3.4, depending on the case of number fields or function fields.
Definition 3.5.
Let $K$ be a global field with $\textup{char}(K)\neq 2$, and keep the constants $a,b,c,d$ as above. We define the ring $R_{p}^{\sigma}$ for $\sigma\in\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)$ as follows:
(a): If $K$ is a number field, then we define
  • $R_{p}^{(-1,-1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{b,p}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p}^{(1,-1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{ab,p}\cap\Delta_{b,p}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p}^{(-1,1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{ab,p}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p,q}^{(1,1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{ap,q}\cap\Delta_{bp,q}}\mathcal{O}_{\mathfrak{p}}$.
(b): If $K$ is a global function field with $\textup{char}(K)\neq 2$, then we define
  • $R_{p}^{(-1,-1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{b,p}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p}^{(1,-1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{ab,p}\cap\Delta_{b,p}\cap\Delta_{a,cp}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p}^{(-1,1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{a,p}\cap\Delta_{ab,p}\cap\Delta_{b,dp}}\mathcal{O}_{\mathfrak{p}}$,
  • $R_{p,q}^{(1,1)}:=\bigcap\limits_{\mathfrak{p}\in\Delta_{ap,q}\cap\Delta_{bp,q}}\mathcal{O}_{\mathfrak{p}}$.
For $K=\mathbb{Q}$, the above definition of $R_{p}^{\sigma}$ and $R_{p,q}^{(1,1)}$ is given as Definition 7 in [Koe16]. For the generalization to number fields, it is Definition 3.10 of [Par13]. For global fields, it is Definition 3.12 of [EM18].
Proposition 3.6.
Let $\Delta$ be a finite subset of $S_{K}^{f}$. Recall that $J(R)$ is the Jacobson radical of $R$. If $R=\bigcap\limits_{\mathfrak{p}\in\Delta}\mathcal{O}_{\mathfrak{p}}$, then $J(R)=\bigcap\limits_{\mathfrak{p}\in\Delta}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}.$
Proof.
Let $x\in R$, the inverse $x^{-1}\in K$ lies in $R$ if and only if $v_{\mathfrak{p}}(x^{-1})\geq 0,\forall\mathfrak{p}\in\Delta$. Combining this with $x\in R$ we get $v_{\mathfrak{p}}(x)=0,\forall\mathfrak{p}\in\Delta$. Hence the set of invertible elements in $R$ is
$\{x\in R:v_{\mathfrak{p}}(x)=0,\forall\mathfrak{p}\in\Delta\}.$
Suppose that $x\in\bigcap\limits_{\mathfrak{p}\in\Delta}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$, then $v_{\mathfrak{p}}(1+yx)=0,\forall y\in R,\forall\mathfrak{p}\in\Delta$, hence $x\in J(R)$.
Suppose that $x\in J(R)$, we claim that $v_{\mathfrak{p}}(x)>0$ for all $\mathfrak{p}\in\Delta$, hence $x\in\bigcap\limits_{\mathfrak{p}\in\Delta}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}.$
If there exists $\mathfrak{p}^{\prime}\in\Delta$ such that $v_{\mathfrak{p}^{\prime}}(x)=0$, then $x^{-1}\in\mathcal{O}_{\mathfrak{p}^{\prime}}$, then by Chinese remainder theorem there exists $y\in R$ with $v_{\mathfrak{p}^{\prime}}(y-x^{-1})\geq 1$. So $v_{\mathfrak{p}^{\prime}}(xy-1)\geq 1$. Hence 1xy1-xy is not invertible in $R$, which contradicts to $x\in J(R)$.
In conclusion we have $J(R)=\bigcap\limits_{\mathfrak{p}\in\Delta}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}.$
Proposition 3.7.
Let $\Delta_{1},\Delta_{2}$ be two finite subsets of $S_{K}^{f}$, then
(1) $\bigcap\limits_{\mathfrak{p}\in\Delta_{1}}\mathcal{O}_{\mathfrak{p}}+\bigcap\limits_{\mathfrak{p}\in\Delta_{2}}\mathcal{O}_{\mathfrak{p}}=\bigcap\limits_{\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}}\mathcal{O}_{\mathfrak{p}}$,
(2) $\bigcap\limits_{\mathfrak{p}\in\Delta_{1}}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}+\bigcap\limits_{\mathfrak{p}\in\Delta_{2}}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}=\bigcap\limits_{\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}}\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$.
Proof.
For both (1) and (2), the left hand sides and obviously contained in the right hand sides. For the other direction, let $t=0$ or $1$. If $x\in K$ is an element such that $v_{\mathfrak{p}}(x)\geq t$ for $\mathfrak{p}\in\Delta_{1}\cap\Delta_{2}$, then by Chinese remainder theorem, there exists $x_{1}\in K$ such that $v_{\mathfrak{p}}(x_{1})\geq t,\forall\mathfrak{p}\in\Delta_{1}$ and $v_{\mathfrak{p}}(x_{1}-x)\geq t,\forall\mathfrak{p}\in\Delta_{2}\backslash\Delta_{1}$. Taking $x_{2}=x-x_{1}$, we have $x=x_{1}+x_{2}$ as the form of left hand sides. ∎
The following definition is a summary of Definition 3.21 and Definition 3.24 in [Par13], which are also written in [EM18] as Definition 3.13 and Definition 3.16.
Definition 3.8.
For each $\sigma\in\textup{Gal}(L/K)$, let
$\Phi_{\sigma}=\{p\in K^{\times}:(p)\in I_{\mathfrak{m}},\psi_{L/K}((p))=\sigma,\textup{ and }\mathbb{P}(p)\subset\mathbb{P}^{(1,1)}\cup\mathbb{P}^{\sigma}\};$
$\widetilde{\Phi_{\sigma}}:={K^{\times}}^{2}\cdot\Phi_{\sigma};$
$\Psi_{K}:=\Big\{(p,q)\in\widetilde{\Phi_{(1,1)}}\times\widetilde{\Phi_{(-1,-1)}}|\prod_{\mathfrak{p}|\mathfrak{m}}(ap,q)_{\mathfrak{p}}=-1\textup{ and }p\in a\cdot{K^{\times}}^{2}\cdot(1+J(R_{q}^{(-1,-1)}))\Big\}.$
For our convenience to use, we conclude some results in [EM18] and [Par13] as follows.
Proposition 3.9.
Let $K$ be a global field and keep notations above. The following holds:
  • (1)
    The sets $\Phi_{\sigma},\Psi_{K},R_{p}^{\sigma},J(R_{p}^{\sigma}),R_{p,q}^{(1,1)},J(R_{p,q}^{(1,1)})$ are all diophantine sets over $K$.
  • (2)
    For any $p\in\Phi_{\sigma}$ and $\sigma\in\textup{Gal}(L/K)$ with $\sigma\neq(1,1)$, $\mathbb{P}^{\sigma}(p)$ is nonempty.
  • (3)
    Let $\sigma\in\textup{Gal}(L/K)$ with $\sigma\neq(1,1)$, and let $\mathfrak{p}\nmid\mathfrak{m}$ be a prime of $K$ satisfying $\psi_{L/K}(\mathfrak{p})=\sigma$. Then there is an element $p\in\Phi_{\sigma}$ such that $\mathbb{P}^{\sigma}(p)=\{\mathfrak{p}\}$.
  • (4)
    For $(p,q)\in\Psi_{K}$, we have $\emptyset\neq\Delta_{ap,q}\cap\Delta_{bp,q}\subset I_{\mathfrak{m}}$.
  • (5)
    For each prime ideal $\mathfrak{p}_{0}$ satisfying $\mathfrak{p}_{0}\nmid\mathfrak{m}$ and $\psi_{L/K}(\mathfrak{p}_{0})=(1,1)$, there exists $(p,q)\in\Psi_{K}$ such that $\Delta_{ap,q}\cap\Delta_{bp,q}=\{\mathfrak{p}_{0}\}$.
Proof.
This is a conclusion of Lemma 3.14 and Lemma 3.17 in [EM18] for global function fields, and of Lemma 3.22 and Lemma 3.25 in [Par13] for number fields. ∎
Remark 3.10.
In fact, the sets $R_{p}^{\sigma},J(R_{p}^{\sigma}),R_{p,q}^{(1,1)},J(R_{p,q}^{(1,1)})$ are not only diophantine, but also in a uniform way rely on $p,q$. In other words, the sets
$\{(p,x)\in K:x\in R_{p}^{\sigma}\}$
$\{(p,x):x\in J(R_{p}^{\sigma})\}$
$\{(x,p,q):x\in R_{p,q}^{(1,1)}\}$
$\{(x,p,q):x\in J(R_{p,q}^{(1,1)})\}$
are diophantine. For these rings, the uniformity comes from the uniformity of $T_{a,b}$ (see Remark 3.2), Lemma 3.1 and Proposition 3.7(1). For these Jacobson ideals, the uniformity comes from Lemma 3.17 of [Par13] and Lemma 3.7(2).
Besides these properties above, the following lemma is implicit in [Par13] and [EM18] indeed, although not written as a lemma or theorem.
Lemma 3.11.
In (5) of Proposition 3.9, we can require that $v_{\mathfrak{p}_{0}}(p)=1,v_{\mathfrak{p}_{0}}(q)=0$ and $q$ is not a square modulo $\mathfrak{p}_{0}$.
Proof.
For the case of number fields, the proof of Lemma 3.25 (c) in [Par13] is made by a specific construction of $(p,q)$. When choosing $p,q\in K$, the condition (B) for $q$ is that the Legendre symbol $(\frac{q}{\mathfrak{p}_{0}})$ is equal to 1-1. This shows that $v_{\mathfrak{p}_{0}}(q)=0$ and $q$ is not a square modulo $\mathfrak{p}_{0}$. The condition (4) for $p$ is $(q,p)_{\mathfrak{p}_{0}}=-1$. By the calculation of Hilbert symbol, we have
$-1=(q,p)_{\mathfrak{p}_{0}}$ $=(-1)^{v_{\mathfrak{p}_{0}}(q)v_{\mathfrak{p}_{0}}(p){\frac{|\mathbb{F}_{\mathfrak{p}_{0}}|-1}{2}}}\big(\frac{q}{\mathfrak{p}_{0}}\big)^{v_{\mathfrak{p}_{0}}(p)}\big(\frac{p}{\mathfrak{p}_{0}}\big)^{v_{\mathfrak{p}_{0}}(q)}$
$=\big(\frac{q}{\mathfrak{p}_{0}}\big)^{v_{\mathfrak{p}_{0}}(p)},$
which implies that $v_{\mathfrak{p}_{0}}(p)$ is odd, which can be modified to $1$ after a free multiplication of a square in $K^{\times}$.
For the case of global function fields, it is in the proof of Lemma 3.17 (3) of [EM18]. In the construction, we also have the property that the Legendre symbol $(\frac{q}{\mathfrak{p}_{0}})$ is equal to 1-1. Since $\mathfrak{p}_{0}\in\Delta_{ap,q}$, we have $(ap,q)_{\mathfrak{p}_{0}}=-1$. By a calculation of the Hilbert symbol, and the fact that $\mathfrak{p}\nmid a$, we conclude $v_{\mathfrak{p}_{0}}(p)$ is odd, as stated in the third-to-last line of the proof of Theorem 1.3 in [EM18]. ∎
3.2. Diagonalization of quadratic forms
Let $f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$ be a quadratic form of $m$ variables. We define the matrix $A=(A_{ij})$ by $A_{ii}=a_{ii}$ and $A_{ij}=\frac{a_{ij}+a_{ji}}{2}$ for $i\neq j$. Let $n$ be an integer from 0 to $m$.
$f$ is of rank $n$ if and only if there exists $(b_{i})_{1\leq i\leq n}$, in which $b_{i}$ are non-zero elements in $K$, and an invertible matrix $C:=(C_{rs})\in M_{m\times m}(K)$ such that $C^{t}AC=B$, where $B$ is a diagonal matrix of rank $n$: $\text{diag}(b_{1},...,b_{n},0,...,0)$.
For $2m^{2}$ variables $t_{ij},u_{rs}$, where $1\leq i,j,r,s\leq m$, we define $m^{2}$ polynomials $h_{kl},1\leq k,l\leq m$ as follows:
First we define the matrices of variables $T$ and $U$ by $T:=(t_{ij})_{1\leq i,j\leq m}$, $U:=(u_{rs})_{1\leq r,s\leq m}$. Then $U^{t}TU$ is a matrix whose entries are polynomials with variables $u_{rs}$ and $t_{ij}$. The polynomials $h_{kl}(t_{ij},u_{rs})$ are defined by the $(i,j)$-th entries of $U^{t}TU$. Moreover, the equation $C^{t}AC=B$ holds if and only if
  • $h_{k,l}(A_{ij},C_{ij})=0$ if $k\neq l$,
  • $b_{k}=h_{k,k}(A_{ij},C_{rs})$, if $1\leq k\leq n$
  • $h_{k,k}(A_{ij},C_{rs})=0$, if $n+1\leq k\leq m$.
In conclusion, we have the following property.
Proposition 3.12.
Let $f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$ be a quadratic form in $m$ variables. Then there exist $m^{2}$ many polynomials $(h_{kl})_{1\leq k,l\leq m}\in\mathbb{Z}[t_{ij},u_{rs}]$ , depending only on the number of variables $m$, such that the following are equivalent:
(1) $f$ does not represent 0;
(2) There exist $b_{1},\dots,b_{m}\in K^{\times}$ such that the diagonal form $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ does not represent 0, and the system of equations
$\begin{cases}h_{kl}(A_{ij},u_{rs})&=0,\quad 1\leq k\neq l\leq m\\ h_{kk}(A_{ij},u_{rs})&=b_{k},\quad 1\leq k\leq m\end{cases}$
admits a solution over $K$ in the variables $u_{rs}$.
Proof.
Let the polynomials $h_{kl}$ be the polynomials defined as above. Those polynomials only depends on $m$. If $f$ is of rank $m$, then the condition that there exist $b_{1},...,b_{m}\in K^{\times}$ such that the equation system admits a solution means exactly that $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ is a diagonalization of $f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$. Then $f$ does not represent 0 is equivalent to $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ does not represent 0. If the rank of $f$ is less than $m$, then $f$ always represents 0. ∎
Under the condition that our main theorem (Theorem 1.2) holds, Corollary 1.4 can be obtained by using the diagonalization process (Proposition 3.12) above. Now we derive Corollary 1.4 from Theorem 1.2.
Proof.
Let $f=\sum_{1\leq i,j\leq m}a_{ij}x_{i}x_{j}$. By Proposition 3.12, the condition for $a_{ij}$ such that $f$ does not represents 0 is equivalent to the following condition:
$\exists b_{1},...,b_{m}\in K^{\times}$ and $C_{rs}\in K$ for $1\leq r,s\leq m$ such that
  • (1)
    $h_{k}(A_{ij};C_{rs})=b_{k}$ for all $1\leq k\leq m$,
  • (2)
    $\sum_{i=1}^{m}b_{i}x_{i}^{2}$ does not represent 0.
Then one can change the condition (2) to an equivalent diophantine condition by Theorem 1.2. ∎
4. Proof of the main results
4.1. Dealing with a Single Prime
Our strategy is to partition the set of primes into finitely many classes and handle the primes in each class in a uniform manner. As a first step, we need to consider the case of a single prime.
The following lemma, which is Lemma 4.4 in [EM18], is repeated here for completeness, as it constitutes a key ingredient in our proof.
Lemma 4.1.
Let $\mathfrak{p}$ be a finite prime of $K$ with $|\mathbb{F}_{\mathfrak{p}}|$ odd. Fix $p,s\in K^{\times}$ so that $v_{\mathfrak{p}}(p)=1$ and $v_{\mathfrak{p}}(s)=0$ so that $\textup{red}_{\mathfrak{p}}(s)$ is not a square in the residue field of $\mathfrak{p}$. Then for $x,y\in K^{\times}$, $(x,y)_{\mathfrak{p}}=-1$ if and only if
$((x\in p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})\wedge(y\textup{ or }-xy\in s\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})))$
$\vee((y\in p\cdot{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})\wedge(x\textup{ or }-xy\in s\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}))).$
Besides, we will also need the following easy observation to obtain an equivalent form of Lemma 4.1.
Lemma 4.2.
Let $\mathfrak{p}$ be a finite prime of $K$ with $|\mathbb{F_{\mathfrak{p}}}|$ odd. Fix $p,s\in K^{\times}$ so that $v_{\mathfrak{p}}(p)=1$ and $v_{\mathfrak{p}}(s)=0$ so that $\textup{red}_{\mathfrak{p}}(s)$ is not a square in the residue field of $\mathfrak{p}$. Then for $x,y\in K^{\times}$,
  • (1)
    $v_{\mathfrak{p}}(x)$ is even if and only if $x\in{K^{\times}}^{2}\cdot{\mathcal{O}_{\mathfrak{p}}^{\times}}$,
  • (2)
    $v_{\mathfrak{p}}(x)$ is odd if and only if $x\in p\cdot{K^{\times}}^{2}\cdot{\mathcal{O}_{\mathfrak{p}}^{\times}}$,
  • (3)
    $x$ is a square modulo $\mathfrak{p}$ if and only if $x\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$,
  • (4)
    $x$ is not a square modulo $\mathfrak{p}$ if and only if $x\in s\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})$.
Proof.
(1): The valuation $v_{\mathfrak{p}}(x)$ is even if and only if the valuation can become 0 after a multiplication of a square, which is equivalent to $x\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}$.
(2) $v_{\mathfrak{p}}(x)$ is odd if and only if $v_{\mathfrak{p}}(x/p)$ is even.
(3) This is a restatement of $x$ is a square modulo $\mathfrak{p}$.
(4) This is due to the fact that the product of two quadratic nonresidue is a quadratic residue. ∎
Lemma 4.3.
With the same notations as in Lemma 4.1, for $x,y\in K^{\times}$, $(x,y)_{\mathfrak{p}}=1$ if and only if
$((x\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})\vee(y\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}})))$
$\wedge((y\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times})\vee(x\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}))).$
Proof.
From Lemma 4.2 we can see that this lemma is a equivalent form of Lemma 4.1. ∎
Theorem 4.4.
Let $\mathfrak{p}$ be a finite or real infinite prime of $K$. The set
$\{(a_{1},a_{2},a_{3},a_{4})\in K^{\times}\times K^{\times}\times K^{\times}\times K^{\times}:f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2}\textup{ is anisotropic over }K_{\mathfrak{p}}\}$
is diophantine over $K$.
Proof.
If $\mathfrak{p}\in S^{f}(K)$, then $f$ is isotropic over $K_{\mathfrak{p}}$ if and only if $a_{1}a_{2}a_{3}a_{4}\in K_{\mathfrak{p}}^{\times}$ and $(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}$. So we just need to verify the following three kinds of sets are diophantine:
  • (1)
    $\{x\in K^{\times}:x\in{K_{\mathfrak{p}}^{\times}}^{2}\}$
  • (2)
    $\{(x,y)\in K^{\times}\times K^{\times}:(x,y)_{\mathfrak{p}}=1\}$
  • (2)
    $\{(x,y)\in K^{\times}\times K^{\times}:(x,y)_{\mathfrak{p}}=-1\}$.
If $|\mathbb{F}_{\mathfrak{p}}|$ is odd, then it follows from Lemma 4.1 and Lemma 4.3 that (2) and (3) are diophantine sets since $\mathcal{O}_{\mathfrak{p}}$ and $\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$ are diophantine over $K$.
If $K$ is a number field and $\mathfrak{p}|2$ then it was stated in the proof of Theorem 4.5 of [EM18] that $K_{\mathfrak{p}}^{\times}:{K_{\mathfrak{p}}^{\times}}^{2}]$ is finite (see also 63:9 in [O’M00]). Let $s_{1},...s_{n}\in K$ be a set of representatives for ${K_{\mathfrak{p}}}^{\times}/{{K_{\mathfrak{p}}}^{\times}}^{2}$ and $S_{i}:=s_{j}\cdot{K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}})$ where $e$ is the absolute ramification index. Then $x\in{K_{\mathfrak{p}}^{\times}}^{2}$ if and only if
$x\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}\vee x\in{K^{\times}}^{2}\cdot(1+\mathfrak{p}^{2e+1}\mathcal{O}_{\mathfrak{p}}).$
$(x,y)_{\mathfrak{p}}=1$ if and only if
$(x,y)\in\bigcup_{i,j:(s_{i},s_{j})_{\mathfrak{p}}=1}S_{i}\times S_{j}.$
$(x,y)_{\mathfrak{p}}=-1$ if and only if
$(x,y)\in\bigcup_{i,j:(s_{i},s_{j})_{\mathfrak{p}}=-1}S_{i}\times S_{j}.$
For a real infinite prime $\mathfrak{p}$, Let $\omega:K\to\mathbb{R}$ be the corresponding embedding of $K$ in to $K_{\mathfrak{p}}=\mathbb{R}$. The Lemma 4.1 of [EM18] say that the set $\{x\in K^{\times}:\omega(x)>0\}$ is diophantine over $K$. Then $f$ is anisotropic if and only if
$(\omega(a_{1})>0\wedge\omega(a_{2})>0\wedge\omega(a_{3})>0\wedge\omega(a_{4})>0)\vee(\omega(a_{1})<0\wedge\omega(a_{2})<0\wedge\omega(a_{3})<0\wedge\omega(a_{4})<0).$
Therefore, these sets are all diophantine sets.
4.2. The proof of main theorem
Now we can prove our main theorem: Theorem 1.2.
Lemma 4.5.
The set
$\{(a_{1},a_{2},a_{3},a_{4})\in K^{\times}\times K^{\times}\times K^{\times}\times K^{\times}:f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2}\textup{ is anisotropic over some }\mathfrak{p}\in S_{K}\}$
is a diophantine set.
Proof.
For $\sigma\in\textup{Gal}(K(\sqrt{a},\sqrt{b})/K)$ with $\sigma\neq(1,1)$, let $s_{\sigma}=a$ if $\sigma=(-1,\pm 1)$ and $s_{\sigma}=b$ if $\sigma=(1,-1)$.
We claim that $f\textup{ is anisotropic over some }\mathfrak{p}\in S_{K}$ if and only if one of the following conditions holds:
  • $\exists\mathfrak{p}|\mathfrak{m}$ such that $f$ is anisotropic over $K_{\mathfrak{p}}$.
  • $\bigvee_{\sigma\neq(1,1)}\quad\exists p\in\Phi_{\sigma}$ such that $(\dagger_{a_{1},a_{2}}^{+}\wedge\dagger_{-a_{3},-a_{4}}^{-})\vee(\dagger_{a_{1},a_{2}}^{-}\wedge\dagger_{-a_{3},-a_{4}}^{+})\wedge\mathfrak{e}$
    where the formula $\mathfrak{e}$ is $a_{1}a_{2}a_{3}a_{4}\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times}\wedge a_{1}a_{2}a_{3}a_{4}\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))$.
    and a formula $\dagger_{x,y}^{-}$ is defined by
    $((x\in p\cdot{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\wedge(y\textup{ or }-xy\in s_{\sigma}\cdot{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
    $\vee((y\in p\cdot{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\wedge(x\textup{ or }-yx\in s_{\sigma}\cdot{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
    and a formula $\dagger_{x,y}^{+}$ is defined by
    $((x\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\vee(y\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
    $\wedge((y\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\vee(x\textup{ and }-yx\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
  • $\exists(p,q)\in\Psi_{K}$ such that $q\in(R_{p,q}^{(1,1)})^{\times}$ and $(\ddagger_{a_{1},a_{2}}^{+}\wedge\ddagger_{-a_{3},-a_{4}}^{-})\vee(\ddagger_{a_{1},a_{2}}^{-}\wedge\ddagger_{-a_{3},-a_{4}}^{+})\wedge\mathfrak{e}$,
    where a formula $\ddagger_{x,y}^{-}$ is defined by
    $((x\in p\cdot{K^{\times}}^{2}\cdot(R_{p,q}^{(1,1)})^{\times})\wedge(y\textup{ or }-xy\in q\cdot{K^{\times}}^{2}\cdot(1+J(R_{p,q}^{(1,1)}))))$
    $\vee((y\in p\cdot{K^{\times}}^{2}\cdot(R_{p,q}^{(1,1)})^{\times})\wedge(x\textup{ or }-yx\in q\cdot{K^{\times}}^{2}\cdot(1+J(R_{p,q}^{(1,1)}))))$
    and a formula $\ddagger_{x,y}^{+}$ is defined by
    $((x\in{K^{\times}}^{2}\cdot(R_{p,q}^{(1,1)})^{\times})\vee(y\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+J(R_{p,q}^{(1,1)}))))$
    $\wedge((y\in{K^{\times}}^{2}\cdot(R_{p,q}^{(1,1)})^{\times})\vee(x\textup{ and }-yx\in{K^{\times}}^{2}\cdot(1+J(R_{p,q}^{(1,1)}))))$
For the first situation, it gives a diophantine set since there are at most finitely many primes in $\mathfrak{m}$, as well as Theorem 4.4.
The second and third situation also give diophantine sets by Proposition 3.9(1) as well as Remark 3.10.
As a conclusion, the set
$\{(a_{1},a_{2},a_{3},a_{4})\in K^{\times}\times K^{\times}\times K^{\times}\times K^{\times}:f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{2}+a_{4}x_{4}^{2}\textup{ is anisotropic over some }\mathfrak{p}\in S_{K}\}$
is a diophantine set, as a union of that three conditions.
It is left for us to show that the claim holds.
[…]
$(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}\textup{ and }a_{1}a_{2}a_{3}a_{4}\in(K_{p}^{\times})^{2}.$
So it is sufficient to prove the following two equivalent relations between the statements in our claim and this condition for different types of $\mathfrak{p}$.
  • For $\sigma\neq(1,1)$, the sentence
    $\exists p\in\Phi_{\sigma}\textup{ such that }\dagger\wedge\mathfrak{e}$
    is equivalent to the condition that there exists a prime $\mathfrak{p}$ with $\psi_{L/K}(\mathfrak{p})=\sigma$ such that $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$ and $a_{1}a_{2}a_{3}a_{4}\in(K_{\mathfrak{p}}^{\times})^{2}$.
  • The sentence
    $\exists(p,q)\in\Psi_{K}\textup{ such that }q\in(R_{p,q}^{\sigma})^{\times}\textup{ and }\ddagger\wedge\mathfrak{e}$
    is equivalent to the condition that there exists a prime $\mathfrak{p}$ with $\psi_{L/K}(\mathfrak{p})=(1,1)$ such that $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$ and $a_{1}a_{2}a_{3}a_{4}\in(K_{p}^{\times})^{2}$.
For $\sigma\neq(1,1)$, if there exists $p\in\Phi_{\sigma}$ such that $\dagger$ holds, then by Proposition 3.9(2) we know that $\mathbb{P}^{\sigma}(p)$ is nonempty. For $\mathfrak{p}\in\mathbb{P}^{\sigma}(p)$, recall that the formula $\dagger_{x,y}^{-}$ is
$((x\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\vee(y\textup{ and }-xy\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
$\wedge((y\in{K^{\times}}^{2}\cdot(R_{p}^{\sigma})^{\times})\vee(x\textup{ and }-yx\in{K^{\times}}^{2}\cdot(1+J(R_{p}^{\sigma}))))$
By Lemma 3.4, and the definition of $R_{\mathfrak{p}}^{\sigma}$, we have $R_{p}^{\sigma}\subset\mathcal{O}_{\mathfrak{p}}$. Also by Proposition 3.6 we have $J(R_{p}^{\sigma})\subset\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$. Hence $\dagger_{x,y}^{-}$ implies $(x,y)_{\mathfrak{p}}=-1$ by $v_{\mathfrak{p}}(p)=1$, Lemma 4.1 and the fact that $s_{\sigma}$ is a $\mathfrak{p}$-adic unit which is not a square by the construction of $s_{\sigma}$ and the definition of Artin map. For the same reason, $\dagger_{x,y}^{+}$ implies $(x,y)_{\mathfrak{p}}=1$, by Lemma 4.2. The formula $\mathfrak{e}$ implies
$a_{1}a_{2}a_{3}a_{3}\in{K^{\times}}^{2}\cdot\mathcal{O}_{\mathfrak{p}}^{\times}\cap{K^{\times}}^{2}\cdot(1+\mathfrak{p}\mathcal{O}_{\mathfrak{p}}),$
which means $a_{1}a_{2}a_{3}a_{4}$ is a square in $\mathcal{O}_{\mathfrak{p}}$. So $\dagger\wedge\mathfrak{e}$ will imply $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$.
Conversely, if there exists a prime $\mathfrak{p}\nmid\mathfrak{m}$ with $\psi_{L/K}({\mathfrak{p}})=\sigma$ such that $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$, then by Proposition 3.9(3) we can find $p\in\Phi_{\sigma}$ such that $\mathbb{P}^{\sigma}(p)=\{\mathfrak{p}\}$. In this situation we have $R_{p}^{\sigma}=\mathcal{O}_{\mathfrak{p}}$ for $\sigma\neq(1,1)$. So the formula $\dagger_{x,y}^{\pm}$ can be deduced from $(x,y)_{\mathfrak{p}}=\pm 1$. Hence, the formula $\dagger\wedge\mathfrak{e}$ can be deduced from $(a_{1},a_{2})_{\mathfrak{p}}\neq-(-a_{3},-a_{4})_{\mathfrak{p}}$.
A similar reason can be used to obtain the second equivalent relation, by using Proposition 3.9(4) and (5). Now we specify the process of proof.
Assume $\mathfrak{p}_{0}\in I_{\mathfrak{m}}$ with $\Psi_{L/K}(\mathfrak{p})=(1,1)$. By Lemma 3.11 and Proposition 3.9, we can find $(p,q)\in\Psi_{K}$ with $\Delta_{ap,q}\cap\Delta_{bp,q}=\{\mathfrak{p}_{0}\}$. Moreover, we have $v_{\mathfrak{p}_{0}}(p)=1$, $v_{\mathfrak{p}_{0}}(q)=0$ and $q$ is not a square modulo $\mathfrak{p}$ by Lemma 3.11. Then by Lemma 4.1, Lemma 4.2 and Lemma 4.3, $(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}$ and $a_{1}a_{2}a_{3}a_{4}\in{K_{\mathfrak{p}}^{\times}}^{2}$ implies $\ddagger\wedge\mathfrak{e}$ since $R_{p,q}^{(1,1)}=\mathcal{O}_{\mathfrak{p}}$ and $J(R_{p,q}^{(1,1)})=\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$.
Conversely, if there exists $(p,q)\in\Psi_{K}$, then by Proposition 3.9, $\Delta_{ap,q}\cap\Delta_{bp,q}\neq\emptyset$. For $\mathfrak{p}\in\Delta_{ap,q}\cap\Delta_{bp,q}$ since $q\in({R_{p,q}^{1,1}})^{\times}$ and $(ap,q)_{\mathfrak{p}}=(bp,q)_{\mathfrak{p}}=-1$, we have
$\big(\frac{q}{\mathfrak{p}}\big)^{v_{\mathfrak{p}}(bp)}=\big(\frac{q}{\mathfrak{p}}\big)^{v_{\mathfrak{p}}(ap)}=-1.$
So $q$ is not a square mod $\mathfrak{p}$ and $v_{\mathfrak{p}}(ap),v_{\mathfrak{p}}(bp)$ are odd. Since the support of $a$ and $b$ are disjoint (Lemma 3.4), we get $v_{\mathfrak{p}}(p)$ is odd. Again by Lemma 4.1, Lemma 4.2 and Lemma 4.3, $\ddagger\wedge\mathfrak{e}$ implies $(a_{1},a_{2})_{\mathfrak{p}}=-(-a_{3},-a_{4})_{\mathfrak{p}}$ and $a_{1}a_{2}a_{3}a_{4}\in{K_{\mathfrak{p}}^{\times}}^{2}$ since $R_{p,q}^{(1,1)}\subset\mathcal{O}_{\mathfrak{p}}$ and $J(R_{p,q}^{(1,1)})\subset\mathfrak{p}\mathcal{O}_{\mathfrak{p}}$. ∎
proof of Theorem 1.2: .
We proceed by considering each possible number of variables $m$ separately.
  • $m=1$. For all $a_{1}\neq 0$, we have $f=a_{1}x_{1}^{2}$ is anisotropic, and the set $K^{\times}$ is obviously diophantine.
  • $m=2$. For all $a_{1},a_{2}\neq 0$, $f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$ is isotropic if and only if $a_{1}a_{2}\in{K^{\times}}^{2}$, and the set
    $\{(a_{1},a_{2})\in{K^{\times}}\times K^{\times}:-(a_{1}a_{2})\notin{K^{\times}}^{2}\}$
    is a diophantine set over $K$ by Theorem 1.4 of [EM18].
  • $m=3$. In this case $f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{3}$. Then $a_{1}a_{2}\in{K^{\times}}^{2}$ means exactly $f=0$ has a nontrivial solution for $x_{3}=0$. The condition $f=0$ has a nontrivial solution for $x_{3}\neq 0$ is equivalent to the solubility of the equation $a_{3}=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}$, which means exactly
    $a_{3}=a_{1}(x_{1}^{2}+a_{1}a_{2}(\frac{x_{2}}{a_{1}})^{2})\in a_{1}\textup{N}_{K(\sqrt{a_{1}a_{2}})/K}(K(\sqrt{a_{1}a_{2}})^{\times}).$
    Hence for all $a_{1},a_{2},a_{3}\neq 0$, $f=a_{1}x_{1}^{2}+a_{2}x_{2}^{2}+a_{3}x_{3}^{3}$ is isotropic if and only if $a_{1}a_{2}\in{K^{\times}}^{2}$ or $a_{3}\in a_{1}\cdot\textup{N}_{K(\sqrt{a_{1}a_{2}})/K}(K(\sqrt{a_{1}a_{2}})^{\times})$.
    The sets
    $\{(a_{1},a_{2},a_{3})\in{K^{\times}}\times K^{\times}\times K^{\times}:a_{1}a_{2}\neq K^{2})\}$
    and
    $\{(a_{1},a_{2},a_{3})\in{K^{\times}}\times K^{\times}\times K^{\times}:a_{3}\notin a_{1}\cdot\textup{N}_{K(\sqrt{a_{1}a_{2}})/K}(K(\sqrt{a_{1}a_{2}})^{\times})\}$
    are all diophantine by Theorem 1.3 of [EM18] and Theorem 1.4 of [EM18].
  • $m=4$. This is Lemma 4.5.
  • $m\geq 5$. In this situation $f$ is anisotropic over $K$ if and only if $f$ is anisotropic over some real infinite primes since $f$ is always isotropic over complex primes and finite primes (63:19 of [O’M00]). For a single real infinite prime $\omega$, $f$ is anisotropic over $\omega$ if and only if
    $({\wedge_{i=1}^{m}\omega(a_{i})>0})\vee(\wedge_{i=1}^{m}\omega(-a_{i})>0),$
    which gives a diophantine set by Lemma 4.1 of [EM18]. Since there are only finitely many real primes, the set
    $\{(a_{1},...,a_{m})\in(K^{\times})^{m}:\Sigma_{i}a_{i}x_{i}^{2}\textup{ is anisotropic over }K\}$
    is diophantine.